# Pruning via Merging: Compressing LLMs via Manifold Alignment Based Layer Merging

### v1

@@ 3.3 How Does MKA Combined with Quantization Perform Compared to Pruning Combined with Quantization?
We compare the performance of MKA with the baseline pruning method, ShortGPT Men et al. (2024), on the MMLU dataset using the Llama3-8B, Llama3-70B, Mistral-7B, Llama2-7B, and Llama2-13B models. The results are shown in Table 1.
We can see that the pruned models are able to be further quantized and maintain performance with a higher compression ratio. Notably, at a high compression ratio of around 87.50%, MKA significantly outperforms ShortGPT. Additionally, we achieve excellent results with various quantization methods. For example, on Llama3-8B, at a compression ratio of 85.94%, MKA with SmoothQuant achieves 64.20%, far exceeding ShortGPT with SmoothQuant at 37.66%. Similarly, with the GPTQ quantization method, we achieve 62.98%, surpassing ShortGPT’s 37.00%, and with AWQ, we achieve 61.66%, exceeding ShortGPT’s 35.44%.
@@
We compared the performance of MKA and several other pruning methods on the LLama3-8B model using multiple benchmark datasets at compression ratios of 34.375%, 37.5%, 40.625% and 43.75%. The results are shown in Table 2. From the results, merging can retain performance better compared to pruning. Relative to SparseGPT and ShortGPT, our method can achieve better performance retention, with significant improvements across all datasets. For example, at a compression ratio of 34.375% on the MMLU dataset, our method can outperform ShortGPT by 21.92% and SparseGPT by 20.42%. Similarly, on the HellaSwag dataset, our proposed method can surpass ShortGPT by 18.32% and SparseGPT by 18.32%.
3.5 Are Inter-Layer Knowledge Alignment Similarity Matrices Consistent Across Different Large Models?
We generate layer similarity heatmaps for different models before and after applying MKA. These heatmaps visualize the knowledge alignment and layer merging effects of MKA on various models. Figure 3 presents the similarity heatmaps for Llama-3-8B, Llama-3-70B, Mistral-7B, Llama-2-7B, and Llama-2-13B. We observe that the heatmaps for the later layers of each model exhibit high similarity values, indicating that inter-layer similarity is consistently high in the later layers across different models. This observation supports our layer merging approach. Additionally, when merging the earlier layers, we notice a collapse of the matrix in the final figure, suggesting that earlier layers have a significant influence on later layers. Thus, simple merging operations on the earlier layers of the model are not feasible.
@@ 4.1 Extension to Multimodal and Specialized Models
In addition to its application to large language models, the MKA method shows promising potential for broader adoption across a variety of deep learning architectures. This includes Mixture-of-Experts (MoE) Jiang et al. (2024), and Mamba Gu and Dao (2023); Lieber et al. (2024) models, which can exhibit similar redundancies in their processing layers.The results show in Figure 4. Initial experiments conducted on these diverse architectures have reinforced the viability of our approach. For instance, the similarity matrices generated on jamba Lieber et al. (2024) and Mixtral-8x7B Jiang et al. (2024) applying MKA have shown that Our method can also be generalized to other similar models, but the similarity distributions of jamba and Mixtral-8x7B are slightly different from LLM, and we do not yet know the reason. These experiments further validates the effectiveness of our method across different model types.
@@ 5 Conclusion
In this paper, we have proposed Manifold-Based Knowledge Alignment and Layer Merging Compression (MKA), a novel model compression technique specifically designed to efficiently reduce the size of large language models (LLMs) while maintaining their performance. MKA leverage manifold learning techniques to align knowledge across layers and utilizes the Normalized Pairwise Information Bottleneck (NPIB) measure to identify the most similar layers for merging. By capturing the intricate nonlinear dependencies within LLMs and integrating knowledge from similar layers, MKA achieves remarkable compression ratios without sacrificing model accuracy. We have conducted extensive experiments on a diverse set of benchmark datasets and various state-of-the-art LLMs to rigorously evaluate the effectiveness of MKA in preserving model performance while significantly reducing model size. Our empirical results demonstrate that MKA consistently outperforms existing pruning methods and can achieve even higher compression ratios when combined with quantization techniques.
Limitations
The quality of the manifold learning process in MKA heavily depends on the diversity and representativeness of the layer activations extracted from the input dataset. In our experiments, we used $\sigma$ value of 8 and selected the first question from the 57-question MMLU dataset to extract activations. We observed that the number of questions sampled can significantly impact the manifold learning results. Ensuring the Condition Number remains below 2000 is crucial for maintaining the integrity of the learned manifold representations. If the dataset used for extracting activations does not adequately cover the model’s operational range, the learned manifold representations might fail to capture the true geometric structure of the data.
The current implementation of MKA has been primarily tested on transformer-based architectures. Although we believe that deep neural networks inherently contain redundancies, the applicability and effectiveness of MKA on other neural network architectures, such as convolutional neural networks (CNNs) or recurrent neural networks (RNNs), have not been thoroughly explored. Future research can investigate these architectures to confirm whether MKA can achieve similar compression benefits across different types of neural networks.
Appendix A More Results

### v2

@@ 6.2 How Does MKA Combined with Quantization Perform Compared to Pruning Combined with Quantization?
We compare the performance of MKA with the baseline pruning method, ShortGPT [14], on the MMLU dataset using the Llama2-7B, Llama3-8B, and Mistral-7B models. The results are shown in Table I.
From Table I, we can see that the pruned models are able to be further quantized and maintain performance with a higher compression ratio. Notably, at a high compression ratio of around 87.50%, MKA significantly outperforms ShortGPT. Additionally, we achieve excellent results with various quantization methods. For example, on Llama3-8B, at a compression ratio of 85.94%, MKA with SmoothQuant achieves 64.20%, far exceeding ShortGPT with SmoothQuant at 37.66%. Similarly, with the GPTQ quantization method, we achieve 62.98%, surpassing ShortGPT’s 37.00%, and with AWQ, we achieve 61.66%, exceeding ShortGPT’s 35.44%.
@@
We compared the performance of MKA and several other pruning methods on the LLama3-8B model using multiple benchmark datasets at compression ratios of 34.375%, 37.5%, 40.625% and 43.75%. The results are shown in Table II. From the results, merging can retain performance better compared to pruning. Relative to ShortGPT, our method can achieve better performance retention, with significant improvements across all datasets. For example, at a compression ratio of 34.375% on the MMLU dataset, our method can outperform ShortGPT by 21.92%. Similarly, on the HellaSwag dataset, our proposed method can surpass ShortGPT by 18.32%.
6.4 Are Inter-Layer Knowledge Alignment Similarity Matrices Consistent Across Different Models?
We generate layer similarity heatmaps for different models before applying MKA. These heatmaps visualize the knowledge alignment and layer merging effects of MKA on various models. Figure 3 presents the similarity heatmaps for Llama2-7B, Llama2-13B, Llama-3-8B, Llama3.2-3B, and Mistral-7B models. We observe that the heatmaps for the later layers of each model exhibit high similarity values, indicating that inter-layer similarity is consistently high in the later layers across different models. This observation supports our layer merging approach. Additionally, when merging the earlier layers, we notice a collapse of the matrix in the final figure, suggesting that earlier layers have a significant influence on later layers. Thus, simple merging operations on the earlier layers of the model are not feasible.
6.5 Iterative Nature of MKA
It’s important to note that the MKA method already incorporates an iterative process in its design. For example, when we merge layers 31 and 32, we obtain a fused layer, which is then merged with layer 30 in the next iteration. We have compared this approach with an alternative method where each layer is allowed to merge only once (e.g., merging layers 31 and 32, then 30 and 29 separately). Our experiments on the Llama3-8B model using the MMLU dataset demonstrate that MKA’s iterative approach yields superior performance in terms of minimizing accuracy degradation. The results are presented in Table IV.
@@ 7.1 Extension to Multimodal and Specialized Models
In addition to its application to large language models, the MKA method shows promising potential for broader adoption across a variety of deep learning architectures. This includes Mixture-of-Experts (MoE) [4], and Jamba [58] models, which can exhibit similar redundancies in their processing layers. The results show in Figure 4. Initial experiments conducted on these diverse architectures have reinforced the viability of our approach. For instance, the similarity matrices generated on Mixtral-8x7B [4] and jamba [58] applying MKA have shown that our method can also be generalized to other similar models, but the similarity distributions of Mixtral-8x7B and Jamba are slightly different from LLM, and we do not yet know the reason. These experiments further validates the effectiveness of our method across different model types.
@@ 7.2 Analysis of Similarity Measures
7.3 Further Exploration of the Merging Ratio
To further investigate the impact of $\lambda_{m}$, we conducted experiments using fixed values of $\lambda_{m}$ without considering layer similarities. We test $\lambda_{m}$ values of 0.7, 0.6, 0.5, and 0.4 (assigning higher weight to the lower-numbered layer). The results are shown in Table III.
These results exhibit a somewhat monotonic trend, with performance decreasing as $\lambda_{m}$ moves away from 0.7. However, all performances remain below that of the similarity-based method. This further highlights the importance of adaptive merging ratios based on layer similarities, as in our MKA method, for maintaining model performance during compression.
@@ 7.4 Variations in Accuracy Across Different MMLU Subjects During Layer Merging
TABLE III: Performance comparison of different fixed merging ratios on Llama3-8B using the MMLU dataset.
CR MKA $\lambda_{m}=0.7$ $\lambda_{m}=0.6$ $\lambda_{m}=0.5$ $\lambda_{m}=0.4$
9.38 66.15 66.06(-0.09) 66.05(-0.10) 65.98(-0.17) 65.96(-0.19)
18.75 64.96 63.47(-1.49) 63.32(-1.64) 62.92(-2.04) 62.83(-2.13)
34.38 64.87 61.84(-3.03) 61.52(-3.35) 61.45(-3.42) 61.59(-3.28)
TABLE IV: Comparison of iterative and non-iterative MKA approaches on Llama3-8B using MMLU dataset.
CR MKA (w/o iterative) MKA (w/ iterative)
0.00 66.29 66.29
3.13 66.13 66.13
6.25 61.64 66.26
9.38 47.43 66.15
12.50 35.87 58.08
15.63 47.82 62.94
18.75 42.01 64.96
21.88 42.00 62.92
25.00 39.39 64.28
28.13 40.07 65.01
31.25 30.41 63.99
34.38 26.73 64.87
37.50 25.37 62.05
@@ 8 Conclusion
In this paper, we have proposed Manifold-Based Knowledge Alignment and Layer Merging Compression (MKA), a novel model compression technique specifically designed to efficiently reduce the size of large language models (LLMs) while maintaining their performance. MKA leverages manifold learning techniques to align knowledge across layers and utilizes the Information Bottleneck (IB) measure to identify the most similar layers for merging. By capturing the intricate nonlinear dependencies within LLMs and integrating knowledge from similar layers, MKA achieves remarkable compression ratios without sacrificing model accuracy. We have conducted extensive experiments on a diverse set of benchmark datasets and various state-of-the-art LLMs to rigorously evaluate the effectiveness of MKA in preserving model performance while significantly reducing model size. Our empirical results demonstrate that MKA consistently outperforms existing pruning methods and can achieve even higher compression ratios when combined with quantization techniques.
Acknowledgements
This work is supported by the National Key R&D Program of China (Grant No. 2023YFB3307500). This work is also supported by the National Natural Science Foundation of China (Grant No. 62306087 and No. 62472121), the Natural Science Foundation of Shandong Province (Grant No. ZR2023QF154), Special Funding Program of Shandong Taishan Scholars Project and the Open Project of Anhui Provincial Key Laboratory of Multimodal Cognitive Computation, Anhui University (Grant No. MMC202420). Besides, we sincerely thank the anonymous reviewers for their valuable feedback.